OpenStudy (anonymous):

Find an equation for the ellipse. Center (0, 0); major axis vertical with length 14; length of minor axis is 6

OpenStudy (tkhunny):

Eyeball problem! Just write it down faster than you can think! \(\dfrac{x^{2}}{(6/2)^{2}} + \dfrac{y^{2}}{(14/2)^{2}} = 1\)

OpenStudy (tkhunny):

Do you see where all the parts emerged?